# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v3

@@ Abstract
Quasisymmetry (QS) is a hidden symmetry of the magnetic field strength, $|\mathbf{B}|$, that confines charged particles effectively in a three-dimensional toroidal plasma equilibrium. Here, we show that QS has a deep connection to the underlying symmetry that makes solitons possible. Our approach uncovers a hidden lower dimensionality of $|\mathbf{B}|$ on a magnetic flux surface, which could make stellarator optimization schemes significantly more efficient. Recent numerical breakthroughs (M. Landreman and E. Paul, Phys. Rev. Lett. 128, 035001 (2022)) have yielded configurations with excellent volumetric QS and surprisingly low magnetic shear. Our approach elucidates why the magnetic shear is low in these configurations. Furthermore, given $|\mathbf{B}|$, we deduce an upper bound on the maximum toroidal volume that can be quasisymmetric only from the properties of $|\mathbf{B}|$ and verify it for the Landreman-Paul precise quasiaxisymmetric (QA) stellarator configuration. In the neighborhood of the outermost surface, we show that $|\mathbf{B}|$ approaches the form of the 1-soliton reflectionless potential (I. Gjaja and A. Bhattacharjee, Phys. Rev. Lett. 68, 2413 (1992)). The connection length diverges, indicating the possible presence of an X-point that could potentially be used as basis for a divertor. We present three independent approaches to demonstrate that quasisymmetric $|\mathbf{B}|$ is described by well-known integrable systems such as the Korteweg-de Vries (KdV) equation. The first approach is weakly nonlinear multiscale perturbation theory, which highlights the crucial role that magnetic shear plays in QS. We show that the overdetermined problem of finding quasisymmetric vacuum fields admits solutions for which the rotational transform is highly constrained. We obtain the KdV equation (and, more specifically, Gardner’s equation for certain choices of parameters). Our second approach is non-perturbative and based on ensuring single-valuedness of $|\bm{B}|$, which directly leads to its Painlevé property shared by the KdV equation. Our third approach uses machine learning, trained on a large dataset of numerically optimized quasisymmetric stellarators. We robustly recover the KdV (and Gardner’s) equation from the data.
@@ I Introduction
In plasma physics, a hidden symmetry called quasisymmetry (QS)(Boozer, 1983; Helander, 2014), forces the magnetic field strength, $B=|\mathbf{B}|$, to be symmetric (Boozer, 1983; Rodriguez, 2022), whereas $\bm{B}$ itself depends on all three spatial coordinates, and is thus said to be three-dimensional (3D). The similarity of this symmetry with particle relabelling symmetries (Padhye and Morrison, 1996a; Webb, 2018) is known (Rodriguez, 2022; Weitzner and Sengupta, 2021) but yet to be fully explored. The hidden symmetry in $B$ makes the bounce action (second adiabatic invariant), $J_{\|}$, independent of the field-line label (Helander, 2014), which guarantees excellent particle confinement in a torus. The 3D nature of B allows stellarators to avoid intrinsic difficulties (such as powerful transients) associated with large plasma currents in axisymmetric configurations such as tokamaks.
@@
It is currently unknown whether exact QS can be achieved exactly in a toroidal volume for a magnetic field that satisfies the constraint of force balance in ideal magnetohydrodynamics (MHD) Freidberg (2014). The problem is of fundamental importance and potentially significant for the design of next-generation stellarator fusion reactors. Analytical results (Garren and Boozer, 1991a, b; Plunk and Helander, 2018; Landreman and Sengupta, 2019) seem to indicate that in the presence of scalar plasma pressure, the resulting 3D nonlinear system of PDEs does not have solutions in a toroidal volume (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2019; Constantin et al., 2021; Rodríguez et al., 2020, 2022). In particular, the near-axis expansion (NAE), which is an asymptotic expansion in the distance from the magnetic axis (Mercier, 1974; Solov’ev and Shafranov, 1970; Bernardin et al., 1986), when applied to QS (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2018), becomes overdetermined beyond the second-order. Moreover, NAEs are typically divergent beyond second-order (L. Fu, private communication). Overdetermined problems do not generally have solutions. Even when they have solutions, they can be subject to restrictive constraints. For example, in the particular case of isodynamic magnetic fields, which satisfy a constraint more restrictive than QS (Helander, 2014; Schief, 2003), it has been shown (Palumbo, 1968; Schief, 2003) that if a solution exists, it must be essentially 2D: either axisymmetric like a tokamak or a helically symmetric straight stellarator.
@@
On the other hand, numerical 3D solutions with a precise level of QS and MHD force balance have been obtained through large-scale numerical optimization, hinting at the possible existence of special and physically interesting toroidal stellarator solutions to the overdetermined problem. However, the numerical search over a typically vast parameter space obscures insight into the quasisymmetric stellarator configuration space that permits precise solutions.
In this work, we demonstrate that QS admits a class of analytic $B$ that is a periodic solution (Novikov et al., 1984; Kamchatnov, 2000) of the KdV equation. We proceed along three independent lines of inquiry to support our result. In the first approach, we formulate the overdetermined problem of QS in a vacuum field in close analogy with the equations that arise in the theory of shallow water waves. Analogously to the water wave theory, we highlight the importance of a traveling wave (TW) frame and the existence of longer length scales in QS. Using a weakly nonlinear multiscale perturbation theory, we show that under some assumptions, the magnetic field strength must satisfy the KdV equation whose solutions satisfy the overdetermined set of equations. We can also show that for small rotational transform, the next order corrections that are needed will lead to Gardner’s equation, a generalization of the KdV equation (Miura et al., 1968; Su and Gardner, 1969; Kamchatnov et al., 2012). As in the isodynamic case (Schief, 2003), we show that the rotational transform is not free but highly constrained, and that for a given rotational transform, a perturbative solution of the overdetermination problem of quasisymmetric vacuum fields can be constructed only for some values of the shear. Although the perturbative approach is necessarily approximate, it provides information not only on quasisymmetric $B$ but also the full $\bm{B}$ over the entire plasma volume.
@@
In the second approach, starting from the basic definitions and properties of QS, and requiring only analyticity and single-valuedness of $B$, we show non-perturbatively the connection to the theory of periodic KdV solitons. This approach is complementary to the first approach and is formally exact. In making the connection to soliton theory, we demonstrate how the field-line label independence of the second adiabatic invariant $J_{\|}$ and other integrals of $B$ in QS are directly related to the infinite hierarchy of conserved quantities of the KdV equation. The exact analytic solution for $B$ is valid in the entire toroidal volume and reduces to the NAE predictions close to the axis.
@@
Our third approach is data-driven. We use PySINDyde Silva et al. (2020); Kaptanoglu et al. (2022) to look for relations between $B$ and its derivatives in the large family of quasi-axisymmetric configurations generated in Buller et al. (2024) and the QUASR database Giuliani (2023). PySINDy employs sparse regression techniques and algorithms to discover governing dynamical systems models from noisy and limited data. We show that over a wide range of configurations and parameters, the KdV equation with the traveling wave form for $B$ holds. As described later in the paper, the only exceptions we found are covered by Gardner’s equation.
@@
The connection with periodic KdV solitons allows us to describe $B$ on a magnetic flux surface with only three functions of the magnetic flux, related to the so-called spectral parameters (Novikov et al., 1984) of the KdV equation. We show that these functions are closely related to the spectrum of $B$ in Boozer coordinates Helander (2014) where the hidden symmetry is manifest, thereby highlighting the low dimensionality of the Boozer spectrum. Furthermore, our analysis shows that the periodicity of $B$ is maintained only when these functions are distinct. Consequently, we can estimate the region of validity of volumetric QS based on where the two spectral parameters intersect (discussed below). Remarkably, near the point of intersection, the connection length (period of $B$) begins to diverge, and $B$ approaches the form of a reflectionless potential (a single soliton that decays at infinity) (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990). The diverging connection length is a characteristic of an X-point that could potentially be the basis for a non-resonant divertor (Mioduszewski et al., 2007; Boozer, 2015; Punjabi and Boozer, 2020).
@@
We demonstrate using both data-driven and traditional numerical methods that our analytic theory can describe well the properties of the field strength on a flux surface of the vacuum equilibrium with precise quasi-axisymmetry (QA)Landreman and Paul (2022) and many other optimized quasisymmetric configurations.
@@ II Basic formulation of quasisymmetry
We begin by reviewing some of the known basic properties of QS. Several equivalent (Helander, 2014; Rodriguez, 2022; Freidberg, 2014; Burby et al., 2020; Rodríguez et al., 2020) mathematical descriptions of QS exist. In this Section, we will briefly discuss the most relevant equations and coordinates to describe QS and point out their inter-relationships.
It can be shown that QS implies the existence of a divergence-free vector field $\bm{u}$ that satisfies the system (Rodríguez et al., 2020)
@@
It follows directly from (1c) and the divergence-free nature of $\bm{u},\bm{B}$ that $\bm{u}\cdot\bm{\nabla}$ and $\bm{B}\cdot\bm{\nabla}$ commute, i.e.,
@@
Here, we have used the usual definition of the commutator $[A_{1},A_{2}]\equiv A_{1}A_{2}-A_{2}A_{1}$ for two operators $A_{1},A_{2}$.
The mathematical structure of (1) strongly suggests that $\bm{u}$ can be associated with a particle relabelling symmetry group (Padhye and Morrison (1996a, b); Salmon (1982, 1988); Webb and Zank (2006); Webb (2018) and references therein) with density replaced by the field strength (Rodriguez, 2022; Weitzner and Sengupta, 2021). Therefore, understanding the hidden symmetry underlying QS might also be beneficial in understanding Casimirs (Hameiri, 2004; Padhye and Morrison, 1996a) in fluid dynamics and ideal MHD.
From equation (1c) we can show that
$\bm{u}=(\bm{B}\times\bm{\nabla}\psi-F(\psi)\bm{B})/B^{2}.$ (3)
@@
which is sometimes called the two-term form of QS Rodríguez et al. (2020).
@@
where $\iota$ is the rotational transform, $N$ is the helicity of QS, $2\pi G_{0}/\mu_{0}$ is the poloidal current outside the surface and $\Phi$ is a toroidal angle. Moreover, the two-term form of QS (equivalently $\bm{u}\cdot\bm{\nabla}B=0$) for vacuum fields leads to a TW solution for $B$ (Sengupta et al., 2021) of the form
@@
Another form of QS is the so-called “triple-product” form (Rodriguez, 2022; Rodríguez et al., 2020)
@@
where $f$ must be a function of $\psi$ and $B$. Note that because $\partial_{\ell}B$ can be both positive and negative, we will assume that $f(B,\psi)$ also depends on $\sigma=\text{sign}(\partial_{\ell}B)$.
@@
We note that there is a degree of freedom in choosing the $(\ell,\alpha)$ pair. By transforming to the coordinates (¯$(\overline{\ell},\alpha)$, where
[…]
@@
Thus, there exists a special TW frame (¯$\overline{\ell},\alpha$) in QS where the field strength $B$ is independent of $\alpha$ (Freidberg, 2014). We show the equivalence of the two TW forms of $B$ given by (7) and (16) in Appendix B.1.
@@
[…]
The field line independence of the second adiabatic invariant $J_{\|}$
[…]
@@
The distinctions between the cases of exact symmetry (axisymmetry or helical symmetry), quasisymmetry and omnigeneity must now be made clear. For the special cases of exact symmetry, the corresponding symmetry vectors are the well-known Killing vectors of Euclidean space (Burby et al., 2020), which implies that $\bm{u}$ and $\bm{B}$ are completely decoupled, and the symmetric form of $\bm{B}$ is determined by $\bm{u}$. Therefore, the condition (1b) can be trivially satisfied. Here, we will focus on the case of 3D QS where $\bm{u}$ cannot be posited a priori. Consequently, the condition (1b) will impose a nontrivial constraint. In particular, unlike the case of exact symmetry, the functional form of $f(B,\psi)$ in the $\partial_{\ell}B$ equation must be constrained. We address this question directly in this work.

### v4

# Periodic Korteweg-de Vries soliton potentials generate quasisymmetric magnetic fields

@@ Abstract
Quasisymmetry (QS) is a hidden symmetry of the magnetic field strength, $B=|{\bm{B}}|$, that effectively confines charged particles in a three-dimensional toroidal plasma equilibrium. Here, we show that QS has a deep connection to the underlying symmetry that makes solitons possible. Our approach uncovers a hidden lower dimensionality of $B$ on a magnetic flux surface, which could make stellarator optimization schemes significantly more efficient. Recent numerical breakthroughs (M. Landreman and E. Paul, Phys. Rev. Lett. 128, 035001 (2022)) have yielded configurations with excellent volumetric QS and surprisingly low magnetic shear. Given $B$, it may be possible to deduce an upper bound on the maximum quasisymmetric toroidal volume which depends only on the properties of $B$. This has been verified for the Landreman-Paul precise quasiaxisymmetric (QA) stellarator configuration. In the neighborhood of the outermost surface, we show that $B$ approaches the form of the 1-soliton reflectionless potential (I. Gjaja and A. Bhattacharjee, Phys. Rev. Lett. 68, 2413 (1992)). The connection length diverges, indicating the possible presence of an X-point or cusp that could potentially be used as a basis for a divertor. We present a non-perturbative approach based on ensuring single-valuedness of $B$, which directly leads to its Painlevé property and the KdV and Gardner’s equations. Finally, we use an approach based on machine learning, trained on a large dataset of numerically optimized quasisymmetric stellarators. We robustly recover the KdV and Gardner’s equations from the data.
@@ I Introduction
In plasma physics, a hidden symmetry called quasisymmetry (QS)(H., 1983; Helander, 2014), forces the time-independent magnetic field strength, $B=|\bm{B}|$, to be symmetric (H., 1983; Rodriguez, 2022), whereas the full vector magnetic field $\bm{B}$ itself depends on all three spatial coordinates, and is thus said to be three-dimensional (3D). The similarity of this symmetry to particle relabelling symmetries (Padhye and Morrison, 1996b; Webb, 2018) is known (Rodriguez, 2022; Weitzner and Sengupta, 2021) but yet to be fully explored.
The hidden symmetry in $B$ makes the bounce action (second adiabatic invariant), $J_{\|}$, independent of the field-line label (Helander, 2014), which guarantees excellent particle confinement in a torus. The 3D nature of B allows stellarators enough flexibility to avoid intrinsic difficulties (such as powerful transients) associated with large plasma currents in axisymmetric configurations such as tokamaks.
@@
It is currently unknown whether exact QS can be achieved in a toroidal volume for a magnetic field that satisfies the constraint of force balance in ideal magnetohydrodynamics (MHD) Freidberg (2014). The problem is of fundamental importance and potentially significant for the design of next-generation stellarator fusion reactors. Analytical results (A. and H., 1991b, a; Plunk and Helander, 2018; Landreman and Sengupta, 2019) seem to indicate that in the presence of scalar plasma pressure, the resulting 3D nonlinear system of PDEs does not have solutions in a toroidal volume (A. and H., 1991b, a; Landreman and Sengupta, 2019; Constantin et al., 2021; Rodríguez,E. et al., 2020, 2022). In particular, the near-axis expansion (NAE), which is an asymptotic expansion in the distance from the magnetic axis (Mercier, 1974; Solov’ev and Shafranov, 1970; P. et al., 1986), when applied to QS (A. and H., 1991b, a; Landreman and Sengupta, 2018), becomes overdetermined beyond the second order. Moreover, NAEs are typically divergent (Fu et al., 2025) beyond second order.
Overdetermined problems do not generally have solutions. Even when they have solutions, they may still be subject to restrictive constraints. For example, in the particular case of isodynamic magnetic fields, which satisfy a constraint more restrictive than QS (Helander, 2014; Schief, 2003), it has been shown (Palumbo, 1968; Schief, 2003) that if a solution exists, it must be essentially 2D: either axisymmetric like a tokamak or a helically symmetric straight stellarator.
@@
On the other hand, numerical 3D toroidal stellarator solutions with a precise level of QS and MHD force balance have been obtained via numerical optimization, suggesting the existence of special, physically interesting toroidal solutions to the overdetermined problem. However, the numerical search over a typically vast parameter space obscures insight into the quasisymmetric configuration space that permits precise solutions. In particular, optimized quasisymmetric configurations tend to form clusters in parameter space (Rodríguez et al., 2023; Giuliani et al., 2024), and may be due to a hidden lower-dimensionality that warrants further analysis.
To construct an analytical model of quasisymmetric vacuum fields in a generic non-symmetric toroidal domain, we must first discuss the fundamental requirements of such a theory. First and foremost, we require global toroidal solutions for $B$ that respect double periodicity (both poloidal and toroidal). We restrict ourselves to the class of smooth magnetic fields that are assumed to be analytic. The analyticity requirement can be rigorously justified near the axis (A. and H., 1991b; Landreman and Sengupta, 2018; Sengupta et al., 2024b). Our analyticity requirement currently excludes interesting classes such as piecewise omnigeneous stellarators (Velasco et al., 2024), which can have sharp changes in the gradients of the field strength.
Second, we require field-line independence of $J_{\|}$, as part of the definition of the properties of QS and omnigeneity. The connection with the KdV equation comes from identifying the field-line label $\alpha$ as time. With this identification, we further require that the adiabatic invariant $J_{\|}$ be an exact (time-independent) invariant. Most adiabatic invariants with analytic time-dependent potentials are not exact. For example, it can be shown that over a long time, adiabatic invariants change by (Landau and Lifshitz, 1960; Henrard, 1993) equal to $C\exp{(-1/\epsilon)}$, where $C$ is a positive $\epsilon$-dependent constant where $\epsilon\ll 1$ is the slowness or the adiabatic parameter. Fortunately, there exist privileged time-dependent potentials for which the coefficient $C$ multiplying the exponentially small term vanishes identically. Consequently, the adiabatic invariants become exact invariants (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990). Since $C$ is proportional to the reflection coefficient in the case of above-barrier reflection from a potential well (Keller and Mu, 1991; Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990), such potentials are also called reflectionless potentials. Reflectionless potentials are intimately connected with solitons and integrable systems (Ablowitz, 2011; Novikov et al., 1984) where they naturally appear in the long wavelength limit of periodic cnoidal waves.
Broadening the search to other integrable periodic nonlinear waves leads us to the question, posed by R.S. MacKay, as to whether Painlevé analysis could determine non-axisymmetric QS equilibrium  [^1] R. MacKay, private communication (Simons team meeting 2019, The Painlevé property (PP) should perhaps also be referred to as the Kovalevskaya-Painlevé property since Kovalevskaya used it to find new cases of integrable tops before Painlevé’s work was published (Michéle Audin, Spinning tops: A course on integrable systems, Vol. 51 (Cambridge University Press, 1999)).. The Painlevé analysis, which involves the study of singularities of differential equations in the complex domain, has yielded fundamental results in several physics problems from quantum field theory to optics (Levi and Winternitz, 2013). Analytic continuation from the real to the complex plane is a helpful technique that physicists often utilize in solving linear problems. For nonlinear systems, Kovaleskaya showed that continuation to the complex plane and analyzing singularities can be equally insightful. Moreover, singularities in the complex plane, even those that are far from the real line, can determine and control the behavior of analytic functions on the real line. An important example of this is the doubly periodic elliptic function (Hille, 1997; Whittaker and Watson, 1920) for which, singularities determine its periodicity on the real line.
In our context, the Painlevé analysis offers a useful methodology for ensuring global analyticity and single-valuedness of the quasisymmetric solutions in the toroidal domain, which is a highly nontrivial problem. The Painlevé property, shared by reflectionless and soliton potentials such as cnoidal waves, guarantees robust exact invariants, global analyticity, and the single-valuedness of $B$ in a toroidal domain. Although it is a priori not obvious that quasisymmetric $B$ belongs to such a restricted class of potentials, the cnoidal solution is entirely consistent with the NAE approach within its limit of applicability.
@@
In this work, we demonstrate that QS admits a class of analytic $B$ that is a periodic solution (Novikov et al., 1984; Kamchatnov, 2000) of the KdV equation. Starting from the basic definitions and properties of QS and requiring only the analyticity and single-valuedness of $B$, we show, non-perturbatively, the connection to the theory of periodic KdV solitons. In making the connection to soliton theory, we demonstrate how the field-line label independence of the second adiabatic invariant $J_{\|}$ and other integrals of $B$ in QS are directly related to the infinite hierarchy of conserved quantities of the KdV equation. The exact analytic solution for $B$ is valid in the entire toroidal volume and reduces to the NAE predictions close to the axis.
@@
Our final approach is data-driven. We use PySINDy de Silva et al. (2020); Kaptanoglu et al. (2022) to look for relations between $B$ and its derivatives in the large family of quasi-axisymmetric configurations generated in Buller et al. (2025) and the QUASR database Giuliani (2023). PySINDy employs sparse regression techniques and algorithms to discover governing dynamical system models from noisy, limited data. We show that, over a wide range of configurations and parameters, the KdV equation with the traveling-wave form for $B$ holds. As described later in the paper, the only exceptions we found are covered by Gardner’s equation.
@@
The connection with periodic KdV solitons allows us to describe $B$ on a magnetic flux surface with only three functions of the magnetic flux, which are related to the so-called spectral parameters (Novikov et al., 1984) of the KdV equation. We show that these functions are closely related to the spectrum of $B$ in Boozer coordinates Helander (2014) where the hidden symmetry is manifest, thereby highlighting the low dimensionality of the Boozer spectrum. Furthermore, our analysis shows that the periodicity of $B$ is maintained only when these functions are distinct. Consequently, it may be possible to estimate the region of validity of volumetric QS based on where the two spectral parameters intersect (discussed below). Remarkably, near the point of intersection, the connection length (period of $B$) begins to diverge, and $B$ approaches the form of a reflectionless potential (a single soliton that decays at infinity) (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990). The diverging connection length is a characteristic of an X-point or sharp ridge that could potentially be the basis for a non-resonant divertor (Mioduszewski et al., 2007; Boozer, 2015; Punjabi and Boozer, 2020).
@@
We demonstrate, using both data-driven and traditional numerical methods, that our analytic theory can describe well the properties of the field strength on a flux surface of the vacuum equilibrium with precise quasi-axisymmetry (QA) Landreman and Paul (2022) and some other optimized quasisymmetric configurations as well.
@@ II Basic formulation of quasisymmetry
We begin by reviewing some of the known basic properties of QS. Several equivalent (Helander, 2014; Rodriguez, 2022; Freidberg, 2014; W. et al., 2020; Rodríguez,E. et al., 2020) mathematical descriptions of QS exist. In this Section, we briefly discuss the most relevant equations and coordinates to describe QS and point out their inter-relationships. Here and throughout the paper, we assume the existence of a set of nested toroidal magnetic surfaces, of the kind illustrated in Fig. 1. These assumptions are standard within magnetic confinement theory. For details about the coordinates, see for example Ref. (D’haeseleer et al., 2012).
Figure 1: An illustration of a toroidal surface with toroidal angle ($\Phi$), poloidal angle ($\theta$) and flux-surface label $\psi$. Figure adapted from Ref. (Imbert-Gérard et al., 2024).
It can be shown that QS implies the existence of a divergence-free vector field $\bm{u}$ that satisfies the system (Rodríguez,E. et al., 2020)
@@
In (1c), $\psi$ can be any flux function (i.e. any single-valued function that satisfies $\vec{B}\cdot\nabla\psi=0$). It follows directly from (1c) and the divergence-free nature of $\bm{u},\bm{B}$ that $\bm{u}\cdot\bm{\nabla}$ and $\bm{B}\cdot\bm{\nabla}$ commute, i.e.,
@@
Here, we have used the usual definition of the commutator $A_{1},A_{2}]\equiv A_{1}A_{2}-A_{2}A_{1}$ for two operators $A_{1},A_{2}$.
The mathematical structure of (1) strongly suggests that $\bm{u}$ can be associated with a particle-relabelling symmetry group (Padhye and Morrison (1996b, a); Salmon (1982, 1988); Webb and Zank (2006); Webb (2018) and references therein) with density replaced by the field strength (Rodriguez, 2022; Weitzner and Sengupta, 2021). Therefore, understanding the hidden symmetry underlying QS might also be beneficial in understanding Casimirs (Hameiri, 2004; Padhye and Morrison, 1996b) in ideal MHD and hydrodynamics.
From equation (1c) it follows that
$\bm{u}=(F(\psi)\bm{B}-\bm{B}\times\bm{\nabla}\psi)/B^{2}.$ (3)
@@
which is sometimes called the two-term form of QS Rodríguez,E. et al. (2020).
@@
where $\iota$ is the rotational transform, $N$ is the helicity of QS, $2\pi G_{0}/\mu_{0}$ is the poloidal current outside the surface and $\Phi$ is a toroidal angle. Moreover, the two-term form of QS (equivalently $\bm{u}\cdot\bm{\nabla}B=0$) for vacuum fields leads to a traveling wave (TW) solution for $B$ (Sengupta et al., 2021) of the form
@@
Another form of QS is the so-called “triple-product” form (Rodriguez, 2022; Rodríguez,E. et al., 2020)
@@
where $f$ is an arbitrary function of $\psi$ and $B$. Note that the relation is only valid locally. For $B$ with only one magnetic well, it suffices to introduce branchs cut at the extrema of $B(\ell)$, which effectively makes $f$ depend on $\sigma=\text{sign}(\partial_{\ell}B)$ as an additional parameter.
@@
We note that there is a degree of freedom in choosing the $(\ell,\alpha)$ pair. By transforming to the coordinates $(\overline{\ell},\alpha)$, where
$\overline{\ell}\equiv\ell-\int H(\psi,\alpha)d\alpha,$ (14)
@@
Thus, there exists a special TW frame ($\overline{\ell},\alpha$) for QS where the field strength $B$ is independent of $\alpha$ (Freidberg, 2014). We show the equivalence of the two TW forms of $B$ given by (7) and (16) in Appendix B.
@@
$\partial_{\alpha}\int_{B\leq B_{b}}d\ell\;F(B;\psi)=0.$ (17)
The field-line independence of the second adiabatic invariant $J_{\|}$
$J_{\|}=\oint\sqrt{\mathcal{E}-B}\;d\ell,$ (18)
@@
The distinctions between the cases of exact symmetry (axisymmetry or helical symmetry in a straight cylinder), QS and omnigeneity must now be made clear. For the special cases of exact symmetry, the equations (1) can be solved by letting $u$ be a Killing vector of Euclidean space (W. et al., 2020), which decouples $\bm{u}$ from $\bm{B}$ such that $\bm{B}$ is determined by $\bm{u}$. In other words, $u$ is given by the symmetry of the underlying space and does not need to be solved for. Therefore, the condition (1b), $\bm{u}\cdot\bm{\nabla}B=0$, can be trivially satisfied. Here, we focus on the case of 3D QS where $\bm{u}$ cannot be posited a priori. Consequently, the condition (1b) imposes a nontrivial constraint. In particular, unlike the case of exact symmetry, the functional form of $f(B,\psi)$ in the $\partial_{\ell}B$ equation must be constrained. We address this question directly in this work.